# Growing Trees and Amoebas’ Replications

### v1

@@ 2.2 Orbit lemma
Fix an amoeba $A=(H,s_{1},\dots,s_{k})$. For a vertex $v\in V(H)$, we define the orbit of $v$, denoted by $\mathop{\mathrm{Orb}}(v)$, as the set of vertices $w\in V(H)$ such that there exists an automorphism of $H$ mapping $v$ to $w$. The completion $[A]$ of amoeba $A$ is defined as
@@
Let $\ell$ be the extension parameter and $A=(H,m)$ be an amoeba. If $\ell\in\{1,2\}$ then $A$ is immortal if and only if its completion is.
For other values of $\ell$, an amoeba is partially immortal if and only if its completion is.
@@
We start from the case $\ell\in\{1,2\}$. Suppose that the completion $A^{\prime}$ of $A$ is mortal, that is, it admits a confining tree $T^{\prime}$. It remains to observe that $T^{\prime}$ confines also $A$, so $A$ is mortal by Lemma 3.
Now suppose by contradiction that $A$ is mortal and $A^{\prime}=(H,m^{\prime})$ is not. Lemma 3 implies that there exists a confining tree $T$ of $A$. Since $T$ does not confine $A^{\prime}$ there exists a simplicial copy of $A^{\prime}$ in $T$, that is, for some vertex $v\in V(H)$ it holds
$d_{T}(v)-d_{H}(v)\leq m^{\prime}(v).$
Since $m^{\prime}(v)=\max_{w\in\mathop{\mathrm{Orb}}(v)}m(w)$, there is a vertex $w\in\mathop{\mathrm{Orb}}(v)$ such that $m(w)=m^{\prime}(v)$. Since $v$ and $w$ are in the same orbit of $H$, there exists a copy of $A$ such that
$d_{T}(v)-d_{H}(v)\leq m(v),$
which contradicts the choice of $T$.
For $\ell>2$, similar arguments show that an amoeba is partially immortal if and only if its completion is. The only difference is to use Theorem 4 instead of Lemma 3. ∎
@@ 3 Case of $1$-extensions
Understanding the mortality of general amoebas seems to be very difficult. In this section we focus on the mortality conditions for amoebas which grow in the simplest of ways – by attaching pendant edges. In our notation such amoeba growth corresponds to the case of $1$-extensions. In [8] authors show that for the case $k(A)=2$, any immortal amoeba is necessarily subcubic. The following subsection generalizes those results to the case of arbitrary $k(A)$.
@@ 3.1 Degree conditions
For the rest of this subsection fix an immortal amoeba $A=(H,m)$. Let $(d(v),m(v))$, $v\in H$, be the sequence of vertex degrees and multiplicities of the roots in $H$. We also define
$\tilde{d}(v)=d(v)+m(v)\qquad\text{and}\qquad m=\max_{v\in V(H)}\tilde{d}(v).$
Fix a $1$-generation sequence $G_{0},G_{1},\dots$. The immediate corollary of minimality of $\ell$-generation is the following claim.
@@
If a vertex $v$ of $G_{i}$ is a root of a copy of the amoeba to be extended and $d_{G_{i+1}}(v)>d_{G_{i}}(v)$ then $d_{G_{i}}(v)<\tilde{d}(v)$.
@@
If $d_{G_{i}}(v)\geq\tilde{d}(v)$ then an extension uses existing edges incident to $v$ due to minimality condition. ∎
@@
$d_{G_{i}}(v)\leq m$ for all $i$ and $v\in V(G_{i})$.
@@
These facts put restrictions on degrees of vertices in $H$. Recall that we add pendant edges. So, initially a degree of a new vertex $v$ is 1. Then it can be increased by extension of a copy having a root at $v$. Thus $\min_{v:m(v)>0}d(v)=1$.
To catch the process of degree increasing, we introduce an auxiliary directed graph $D$. The vertices of $D$ are positive integers up to $m$. A pair $(x,y)$ is an edge in $D$ if $y=d(v)+m(v)$ and $d(v)+m(v)>x\geq d(v)$ for some $v\in V(H)$.
Let $q$ be the maximal integer reachable from 1 in $D$. By definition of $q$, the inequality $q\leq m$ holds. For an immortal amoeba $m=d(v^{*})+m(v^{*})$ should be reachable because the degree $d(v^{*})$ should appear in some new vertex (i.e. a vertex from $V(G_{i})\setminus V(G_{0})$) and all copies containing this vertex at some moment have enough space for the extension. This implies the following theorem.
@@
Let $A=(H,m)$ be an immortal amoeba. Then
  1. 1.
    $q=m$;
  2. 2.
    $d_{H}(v)\leq m$ for all $v\in V(H)$.
@@
Let $x_{1},x_{2},\dots,x_{t}$ be a directed path in $D$. By definition of $D$, $x_{i}=d(v_{i})+m(v_{i})$ and $x_{i-1}\geq d(v_{i})$, for $i>1$. Then
@@
Let $A=(H,m)$ be amoeba. If $\Delta(H)>1+k$, then $A$ is mortal.
@@
Let $A=(H,m)$ be an immortal amoeba such that $k=1$. Then $\Delta(H)\leq 2$.
@@ 3.2 Mortality criteria for caterpillars
In this subsection a $1$-extension of a given amoeba will be referred to simply as extension. Furthermore, throughout this section we assume that the multiplicity function $m$ takes values 00 and $1$ only. A caterpillar $C(d_{1},d_{2},\dots,d_{\ell})$ is a tree constructed as follows. We start with a central path $P_{\ell}=(p_{1},\dots,p_{\ell})$, and afterwards, for each $i\in\{1,\dots,\ell\}$, we add $d_{i}$ pendant vertices to $p_{i}$. In addition, we require $d_{1}=d_{\ell}=0$. A caterpillar is said to be symmetric if $d_{i}=d_{\ell-i+1}$, for all $i$. The family of amoebas $(H,m)$ where $H$ is a caterpillar and $m$ respects the above mentioned properties is denoted by $\mathcal{A}$. In this section we characterize immortal amoebas from $\mathcal{A}$.
@@
It is enough to observe that the central path of the obtained graph corresponds to $P$, possibly with additional vertex on either endpoint of $P$ (also see fig. 5). ∎
We now give necessary and sufficient condition for an amoeba from $\mathcal{A}$ to be shifted ”to the right”. But first we need some definitions.
@@
An amoeba $A\in\mathcal{A}$ is immortal if and only if $[A]$ is slow.
@@
Let $S=(d_{1},d_{2},\dots)$ be the degree sequence of the central path of $[A]$. To prove the ‘if direction’ assume wlog that $[A]$ is slowly decreasing. We define a sequence of graphs $\mathcal{G}=(G_{1},G_{2},\dots)$, starting with $G_{1}=C(S)$. (Informally speaking, $\mathcal{G}$ is a sequence of ”right shifts” of $[A]$.) Every member of this sequence is a caterpillar graph, hence, the vertex notation (vertices $p_{1},\dots$) is well defined.
@@
Due to Lemma 3 of the ‘if direction’ of the statement it is enough to prove that $\mathcal{G}$ is infinite. Assume for contradiction that $\mathcal{G}$ has (a finite) length $k$. By definition $G_{1}=C(S)$ contains $\mathcal{C}_{1}$, hence, $k\geq 2$.
@@
Note that (ii) implies $d_{G_{k}}(p_{k+|S|})=0$ which is impossible. Let us fix an $i$ satisfying (i).
By construction, there exists a copy $H$ of $C(S)$ in $G_{k-1}$ containing $\mathcal{C}_{k-1}$ whose extension gives rise to $G_{k}$. By (i) we have $d_{i}-d_{i-1}<0$, furthermore $d_{i}-d_{i-1}=-1$ since $S$ is slowly decreasing. However, due to the extension of $[A]=(H,m^{\prime})$ in $G_{k-1}$, we are guaranteed that $m^{\prime}(p_{k+i})>0$, and hence $d_{G_{k}}(p_{k+i})-2\geq d_{i}$, which is a contradiction.
@@
    $i$ and $j$ such that $d_{i}-d_{i-1}<-1$ while $d_{j}-d_{j-1}>1$;
@@
    $i$ and $j$ such that $d_{i}-d_{i-1}<-1$, and $d_{j}-d_{j-1}=1$ while $p_{j-1}$ is not a root;
@@
    $i$ and $j$ such that $d_{i}-d_{i-1}=-1$, and $d_{j}-d_{j-1}>1$ while $p_{i}$ is not a root;
@@
To finish the proof, take a copy $A^{\prime}$ of $[A]$ in $G_{2}$ such that its central path contains a pendant vertex $v$ from $G_{1}$ adjacent to $p_{\ell-1}$; the case when it is adjacent to $p_{2}$ is proved similarly. To have enough room for $A^{\prime}$ to grow, we add one pendant edge to $v$ only. Notice that no new copies of $[A]$ appear, thus, $[A]$ is mortal, and due to Lemma 5, so is $A$. ∎
@@ 4 Directions for future work
As announced in Section 2, we conjecture that the three notions of immortality are equivalent.
@@
  • $A$ is immortal.
  • $A$ is strongly immortal.
  • $A$ is partially immortal.
Deciding mortality looks hard (undecidable); perhaps one can model a general Turing machine by amoebas. For this purpose a more general notion of an amoeba colony might be useful. An amoeba colony is a set of ordinary amoebas $\mathcal{C}=\{A_{1},\dots,A_{n}\}$. An $\ell$-growth $G^{\prime}$ of $\mathcal{C}$ in $G$ is an $\ell$-growth of some $A_{i}\in\mathcal{C}$. The remaining definitions are the same as for the case of ordinary amoebas. So, in construction of $\ell$-growth sequence for an amoeba colony, one can choose an alive amoeba to be extended. It gives more freedom and we conjecture that such a process can simulate a universal computation.
@@ Acknowledgements.
Authors are grateful to Prof. Martin Milanič for valuable discussions and observations.
This research was partially prepared within the framework of the HSE University Basic Research Program. This work is supported in part by the Slovenian Research Agency (research program P1-0383 and research projects N1-0160, N1-0209, J1-3003, J1-4008 and J5-4596).

### v2

@@ 2.2 Orbit lemma
Fix an amoeba $A=(H,m)$. For a vertex $v\in V(H)$, we define the orbit of $v$, denoted by $\mathop{\mathrm{Orb}}(v)$, as the set of vertices $w\in V(H)$ such that there exists an automorphism of $H$ mapping $v$ to $w$. The completion $[A]$ of amoeba $A$ is defined as
@@
Let $\ell$ be the extension parameter and $A=(H,m)$ be an amoeba. If $\ell\in\{1,2\}$ then $A$ is mortal if and only if its completion is. For other values of $\ell$, an amoeba is mortal w.r.t all trees if and only if its completion is.
@@
We start from the case $\ell\in\{1,2\}$. Suppose that $[A]$ is mortal, that is, it admits a confining tree $T^{\prime}$. It remains to observe that $T^{\prime}$ confines also $A$, so $A$ is mortal by Lemma 3.
Now suppose that $A$ is mortal. Lemma 3 implies that there exists a confining tree $T$ of $A$. Let $(H,m^{\prime})$ be a copy of $[A]$ in $T$, $v\in V(H)$, and $v^{*}$ be a vertex in $\mathop{\mathrm{Orb}}(v)$ such that $m(v^{*})=m^{\prime}(v)$. By the choice of $v^{*}$, there exists an automorphism of $H$ mapping $v^{*}$ to $v$. The corresponding copy of the amoeba $A=(H,m)$ has an $\ell$-extension inside $T$. It means that $H$ can be extended by $m^{\prime}(v)$ paths of length $\ell$ in $T$ such that the paths have the only common vertex $v$ and each of these paths has the only common vertex $v$ with $H$.
By using such paths at each vertex of $H$, one can construct an $\ell$-extension of any copy of $[A]$ inside $T$. It means that $T$ confines $[A]$.
For $\ell>2$, similar arguments show that an amoeba is not mortal w.r.t some tree if and only if its completion is. The only difference is in applying Theorem 4 instead of Lemma 3. ∎
@@ 3 Case of $1$-extensions
Understanding the mortality of general amoebas seems to be very difficult. In this section we focus on the mortality conditions for amoebas which grow in the simplest of ways – by attaching pendant edges. [^3] In our notation such amoeba growth corresponds to the case of $1$-extensions. In Remark 3 we explained that for this case, immortality is just the negation of mortality. The results of [13] easily imply that any immortal amoeba is subcubic whenever the number of roots is equal to $2$. The following subsection generalizes those results to the case of arbitrary number of roots.
@@ 3.1 Degree conditions
For the rest of this subsection fix an immortal amoeba $A=(H,m)$. Let $(d(v),m(v))$, $v\in H$, be the sequence of vertex degrees and multiplicities of the roots in $H$. We also define $\widetilde{m}=\max_{v\in V(H)}(d(v)+m(v))$.
Fix an infinite $1$-generation sequence $G_{0},G_{1},\dots$. The next claim immediately follows from the definition of the $\ell$-extension.
@@
If a vertex $v$ of $G_{i}$ is a root of a copy of the amoeba to be extended and $d_{G_{i+1}}(v)>d_{G_{i}}(v)$ then $d_{G_{i}}(v)<d_{H}(v)+m(v)$.
@@
If $d_{G_{i}}(v)\geq d_{H}(v)+m(v)$ then an extension uses existing edges incident to $v$ due to the minimality condition. ∎
@@
$d_{G_{i}}(v)\leq\widetilde{m}$ for all $i$ and $v\in V(G_{i})$.
This corollary provides a uniform upper bound on vertex degrees in all graphs $G_{i}$. From this we derive the following.
@@
Proof.
If a vertex $v\in V(G_{i})$ belongs to a copy of the amoeba extending at some step, then a new pendant edge is added at distance at most $|V(H)|$ from $v$ in $G_{i+1}$. Since all degrees of vertices in all graphs in the generation sequence are bounded by $\widetilde{m}$ due to 7, it can repeat finitely many times. ∎
These facts put restrictions on degrees of vertices in $H$. Recall that we add pendant edges. So, initially a degree of a new vertex $v$ is 1. Then it can be increased by extension of a copy having a root at $v$. Thus, $\min_{v:m(v)>0}d(v)=1$.
To catch the process of degree increasing, we introduce an auxiliary directed graph $D_{A}$. The vertices of $D_{A}$ are positive integers up to $\widetilde{m}$. A pair $(x,y)$ is an edge in $D_{A}$ if $y=d(v)+m(v)$ and $d(v)+m(v)>x\geq d(v)$ for some $v\in V(H)$.
Let $q_{A}$ be the maximal integer reachable from 1 in $D_{A}$. By the definition of $q$, the inequality $q\leq\widetilde{m}$ holds. For an immortal amoeba, let $v^{*}$ be such that $\widetilde{m}=d(v^{*})+m(v^{*})$. Note that $\widetilde{m}$ is reachable because $d(v^{*})$ appears in a new vertex (i.e. a vertex from $V(G_{i})\setminus V(G_{0})$) and all copies containing this vertex at some moment have enough space for the extension. This implies the following theorem.
@@
Let $A=(H,m)$ be an immortal amoeba with its auxiliary digraph $D_{A}$. Then $q_{A}=\widetilde{m}$ and $d_{H}(v)\leq\widetilde{m}$ for all $v\in V(H)$.
@@
Let $x_{1},x_{2},\dots,x_{t}$ be a directed path in $D$. By definition of $D$, for all $i\in[1,t]$ there exists $v_{i}\in V(H)$ such that $x_{i}=d(v_{i})+m(v_{i})$ and $x_{i-1}\geq d(v_{i})$, for $i>1$. Then
@@
Let $A=(H,m)$ be an immortal amoeba such that $k=1$. Then $\Delta(H)\leq 2$.
@@
Let $A=(H,m)$ be an amoeba. If $\Delta(H)>k+1$, then $A$ is mortal.
The following generalization of the above corollary remains open.
Conjecture 1.
For any value of the extension parameter $\ell$, and amoeba $A=(H,m)$, let $k$ be the number of roots in $A$. If $\Delta(H)>k+2$ then $A$ is totally confined.
@@ 3.2 Mortality criteria for caterpillars
In this subsection a $1$-extension of a given amoeba will be referred to simply as an extension. Furthermore, throughout this section we assume that the multiplicity function $m$ takes values 00 and $1$ only. A caterpillar $C(d_{1},d_{2},\dots,d_{\ell})$ is a tree constructed as follows. We start with a central path $P_{\ell}=(p_{1},\dots,p_{\ell})$, and afterwards, for each $i\in\{1,\dots,\ell\}$, we add $d_{i}$ pendant vertices to $p_{i}$. In addition, we require $d_{1}=d_{\ell}=0$. A caterpillar is said to be symmetric if $d_{i}=d_{\ell-i+1}$, for all $i$. The family of amoebas $(H,m)$ where $H$ is a caterpillar and $m$ respects the above-mentioned properties is denoted by $\mathcal{A}$. In this section we characterize mortality of amoebas from $\mathcal{A}$.
@@
It is enough to observe that the central path of the obtained graph corresponds to $P$, possibly with an additional vertex on either endpoint of $P$ (also see Fig. 6). ∎
We now give a necessary and sufficient condition for an amoeba from $\mathcal{A}$ to be shiftable ‘to the right’. We start with some definitions.
@@
An amoeba $A\in\mathcal{A}$ is mortal if and only if $[A]$ is not slow.
@@
Let $S=(d_{1},d_{2},\dots)$ be the degree sequence of the central path of $[A]$. To prove the ‘if direction’ assume w.l.o.g. that $[A]$ is slowly decreasing. We define a sequence of graphs $\mathcal{G}=(G_{1},G_{2},\dots)$, starting with $G_{1}=C(S)$. (Informally speaking, $\mathcal{G}$ is a sequence of ‘right shifts’ of $[A]$.) Every member of this sequence is a caterpillar graph, hence, the vertex notation (vertices $p_{1},\dots$) is well defined.
@@
Due to Lemma 3 of the ‘if direction’ of the statement, it is enough to prove that $\mathcal{G}$ is infinite. Assume for contradiction that $\mathcal{G}$ has (a finite) length $k$. By definition $G_{1}=C(S)$ contains $\mathcal{C}_{1}$, hence, $k\geq 2$.
@@
Note that (ii) implies $d_{G_{k}}(p_{k+|S|})=0$, which is impossible. Let us fix an $i$ satisfying (i).
By construction, there exists a copy $H$ of $C(S)$ in $G_{k-1}$ containing $\mathcal{C}_{k-1}$ whose extension gives rise to $G_{k}$. By (i) we have $d_{i}-d_{i-1}<0$; furthermore, $d_{i}-d_{i-1}=-1$, since $S$ is slowly decreasing. However, due to the extension of $[A]=(H,m^{\prime})$ in $G_{k-1}$, we are guaranteed that $m^{\prime}(p_{k+i})>0$, and hence $d_{G_{k}}(p_{k+i})-2\geq d_{i}$, which is a contradiction.
@@
    $i$ and $j$ such that $d_{i}-d_{i-1}<-1$, while $d_{j}-d_{j-1}>1$;
@@
    $i$ and $j$ such that $d_{i}-d_{i-1}<-1$, and $d_{j}-d_{j-1}=1$, while $p_{j-1}$ is not a root;
@@
    $i$ and $j$ such that $d_{i}-d_{i-1}=-1$, and $d_{j}-d_{j-1}>1$, while $p_{i}$ is not a root;
@@
To finish the proof, take a copy $A^{\prime}$ of $[A]$ in $G_{2}$ such that its central path contains a pendant vertex $v$ from $G_{1}$ adjacent to $p_{\ell-1}$; the case when it is adjacent to $p_{2}$ is proved similarly. To have enough room for $A^{\prime}$ to grow, we add one pendant edge to $v$ only. Notice that no new copies of $[A]$ appear, thus $[A]$ is mortal, and due to Lemma 5, so is $A$. ∎
@@ 4 Directions for future work
As announced in Section 2, we conjecture the following.
@@
  • $A$ is mortal.
  • $A$ is not immortal.
  • $A$ is mortal w.r.t. any tree.
Deciding mortality looks hard (undecidable); perhaps one can model a general Turing machine by amoebas. For this purpose a more general notion of an amoeba colony might be useful. An amoeba colony is a set of amoebas $\mathcal{C}=\{A_{1},\dots,A_{n}\}$. An $\ell$-growth $G^{\prime}$ of $\mathcal{C}$ in $G$ is an $\ell$-growth of some $A_{i}\in\mathcal{C}$. The remaining definitions are the same as for the case of amoebas. So, in construction of an $\ell$-growth sequence for an amoeba colony, one can choose an alive amoeba to be extended. It gives more freedom and we conjecture that such a process can simulate a universal computation.
@@ Acknowledgements.
The authors are grateful to Martin Milanič for valuable discussions and observations, Susan D. Cook for English proofreading, and to an anonymous reviewer for many helpful remarks.
This research was partially prepared within the framework of the HSE University Basic Research Program. This work is supported in part by the Slovenian Research Agency (research program P1-0383 and research projects N1-0160, N1-0209, J1-3003, J1-4008 and J5-4596). The work of the third author was partially supported by State Assignment, theme no. FFNG-2024-0003.